40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14918     Entry time: Mon Sep 30 18:20:26 2019
Author: gautam 
Type: Update 
Category: ALS 
Subject: ALS OOL noise - a first look 

Attachment #1 shows a first look at the IR ALS noise after my re-coupling of the IR light into the fiber at EY. 

Measurement configuration: 

  • Each arm length was individually stabilized to the PSL frequency using POX/POY locking.
  • The respective AUX laser frequencies were locked to the arm cavity length using the AUX PDH loops.
  • GTRX ~0.3 (usually I can get ~0.5) and GTRY ~ 0.2 (the mode-matching to the arm cavities is pretty horrible as suggested by the multitude of bullseye modes seen when toggling the shutter).
  • The control signal to the AUX PZT had the DC part offloaded by the slow temperature control servos to the AUX laser crystal temperature.

CDS model changes:

  • The c1lsc model was modified to route the input signals to the Y phase tracker servo from ADC1_2 and ADC1_3 (originally, they were ADC0_20 and ADC0_21).
  • This change was necessary because the DFD output is sent differentially to the ADC1 card in the c1lsc expansion chassis (bypassing the iLIGO whitening and AA electronics, for now just going through an aLIGO AA board with no whitening available yet).
  • I chose to use the differential receiving (as opposed to using the front-panel single ended BNC connectors) as in principle, it is capable of delivering better noise performance.
  • After making the model changes, I compiled and restarted the model. Apart from the missing path issue, the compile/restart went smoothly.

Next steps:

  • Get the easy fixes done (better GTRX, GTRY).
  • Test the noise with POX and POY as the OOL sensors, and the arms controlled using the ALS error signal - this is the relevant metric for how ALS will be used in locking.
  • Noise budget. Need to double-check the DFD output calibration into Hz.
  • For the general interferometer recovery, I think I will push ahead with trying to lock some other configurations like the PRMI (should be easy to recover), DRMI (potentially more difficult to find the right settings), and the FPMI (I'd like to use this config to get an estimate for how much contrast defect we have in the interferometer, but I think it'll be pretty challenging to lock in this configuration).
Attachment 1: ALS_OOL_20190930.pdf  75 kB  Uploaded Mon Sep 30 19:21:44 2019  | Hide | Hide all
ALS_OOL_20190930.pdf
ELOG V3.1.3-